40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1270     Entry time: Tue Feb 3 23:44:44 2009
Author: Kakeru, Peter, Yoichi 
Type: Update 
Category: PSL 
Subject: ISS unstability 

We found that one OP-amp used in ISS servo oscillated in 10 MHz, 100mV.

Moreover, we found another OP-amp had big noise.

We guess that these oscilation or noise cause saturation in high frequency, and they effect to lower frequency to cause 

 Attached files are open loop transfar function of ISS.

The blue points are open loop TF, and the green line is product of TF of ISS servo filter and TF of current shunt TF of servo filter.

This two must be same in principle, but They have difference f<2Hz and f>5kHz.

Attachment 1: TFgain.png  9 kB  Uploaded Wed Feb 4 09:41:43 2009  | Hide | Hide all
TFgain.png
Attachment 2: TFphase.png  12 kB  Uploaded Wed Feb 4 09:42:39 2009  | Hide | Hide all
TFphase.png
ELOG V3.1.3-